# Cooperation Does Matter: Exploring Multi-Order Bilateral Relations for Audio-Visual Segmentation

### 11 Dec 2023 v1

@@ Abstract
Recently, an audio-visual segmentation (AVS) task has been introduced, aiming to group pixels with sounding objects within a given video. This task necessitates a first-ever audio-driven pixel-level understanding of the scene, posing significant challenges. In this paper, we propose an innovative audio-visual transformer framework, termed COMBO, an acronym for COoperation of Multi-order Bilateral relatiOns. For the first time, our framework explores three types of bilateral entanglements within AVS: pixel entanglement, modality entanglement, and temporal entanglement. Regarding pixel entanglement, we employ a Siam-Encoder Module (SEM) that leverages prior knowledge to generate more precise visual features from the foundational model. For modality entanglement, we design a Bilateral-Fusion Module (BFM), enabling COMBO to align corresponding visual and auditory signals bi-directionally. As for temporal entanglement, we introduce an innovative adaptive inter-frame consistency loss according to the inherent rules of temporal. Comprehensive experiments and ablation studies on AVSBench-object (84.7 mIoU on S4, 59.2 mIou on MS3) and AVSBench-semantic (42.1 mIoU on AVSS) datasets demonstrate that COMBO surpasses previous state-of-the-art methods. Code and more results will be publicly available at https://combo-avs.github.io/.
Figure 1: Comparison between the proposed COMBO and existing state-of-the-art methods. Our COMBO is the first work to simultaneously explore multi-order bilateral relations in modality, temporal, and pixel levels.
@@ 1 Introduction
Human visual attention is often hear-guided, i.e., we tend to focus on the object with sounds [4]. For example, when we hear a cat meow, we pay more attention to the cat than other objects due to the strong association between the meow and the cat. Inspired by this potential interaction of auditory and visual signals, the cross-modal studies of vision and hearing have attracted numerous researchers, such as the audio-visual correspondence [2, 1], which only aims to match visual images and audio signals to the same scene, and sound source localization [18, 34, 5] which further seeks to locate the vocal visible regions. However, they have only focused on audio-visual tasks at the image or region levels, lacking pixel-level annotations. Recently, AVSBench [45] integrates audio signals into video segmentation, called Audio-Visual Segmentation (AVS), which comprises two benchmarks: 1) AVSBench-object, which includes single source sound segmentation (S4) and multiple sound sources segmentation (MS3); 2) AVSBench-semantic, which further extends audio-visual semantic segmentation (AVSS) based on AVSBench-object.
Given that audio-visual segmentation is a burgeoning field that spans both audio and visual modalities, it presents a non-trivial task. Generally, when performing AVS, a cross-modal segmentation task involving video, there are mainly three challenges: (1) AVS contains audio and visual modalities, thus demanding explicit alignment of sequential audio features to spatial pixel-level activations; (2) AVS involves temporal information where the state of the current frame is dynamically affected by historical frames, therefore, exploring the correlation between adjacent frames is essential; (3) AVS includes an image segmentation task, compared to 1D audio signals, 2D image signals have more heavier redundant information which is prone to be effected by background noise, thus requiring precise extraction of features from the image. To resolve the issue (1), the prevailing methods [31, 33, 13, 26, 25] employ matrix multiplication and modified cross-attention module to encode pixel-wise audio-visual interaction. Though impressive, these designs ignore the temporal dependence of adjacent frames that have been proven to be important for AVS. To solve this problem, some works [45, 15, 23] partition temporal relations into consideration to explore both issues (1) and (2) simultaneously since AVS is a cross-modal video task. Nevertheless, their approach relies too much on implicit inter-frame relations, leading to inaccurate associations. Regarding issue (3), CSMF [3] leverages frozen large foundation models to extract more pure visual features for AVS. However, it independently tackle the audio and visual signals by naively combining several exsiting foundation models, resulting in sub-optimal performance.
@@
In this work, we explore three potential bilateral entanglements: pixel entanglement, modality entanglement, and temporal entanglement. Pixel entanglement refers to the interdependent relationship between an image and its corresponding mask. Since background noise in the image leads to inaccuracies in the image-to-mask prediction process, it is essential to utilize external masks from the foundation models to entangle the input image to assist the model. Therefore, we construct a Siam-Encoder Module (SEM) as a visual feature extractor to facilitate more precise visual features, which can liberate from the constraints of foundation models than other methods [33, 37, 26]. Besides, as for the alignments of the audio and visual signals, we explore the modality entanglement between audio and visual components to amplify the efficiency of cross-modal matching. Contrary to existing single-fusion methods [45, 13], we believe that the cooperation between the two modalities can produce the positive effect. Inspired by [24], we initially propose a potent and memory-efficient bidirectional audio-visual fusion module called Bilateral-Fusion Module (BFM). The BFM amplifies the spatial awareness of visual features relevant to sounding objects and strengthens the attention of audio signals embodying visual target information. Moreover, the audio-visual tasks contain a solid temporal entanglement. Therefore, we design an adaptive inter-frame consistency loss to better harness this inherent characteristic.
@@
  • We show that COMBO significantly outperforms exsiting state-of-the-art approaches in the challenging AVSBench-object and AVSBench-semantic datasets.
@@ 2 Related Work
2.1 Sound Source Localization (SSL)
Sound source localization (SSL) aims to estimate the position of a sound source in a video sequence, which is the most related task to the audio-visual segmentation task. LVS [5] utilizes a hard-mining strategy and contrastive learning mechanism to discriminate challenging image fragments. DSOL [18] executes class-aware sounding object localization from mixed sound, which initially focuses on learning robust object representations from single-source localization. MSSL [34] localizes multiple sound sources in unconstrained videos without pairwise sound-object annotations. This approach involves the development of a two-stage learning framework, followed by the execution of cross-modal feature alignment. These pioneering works in these areas have significantly inspired our research on audio-visual segmentation.
@@ 2.2 Semantic Segmentation
Semantic segmentation is a fundamental task that requires pixel-level classification. Early researchers take the Fully Convolutional Networks (FCN) [27] as the dominant approach and focus on aggregating long-range context in the feature map. PSPNet [42] performs spatial pyramid pooling at several grid scales. DeepLab [7, 8] utilizes atrous convolutions with different atrous rates. Furthermore, some methods [38, 35, 43] replace traditional convolutional backbone with transformer-based architectures. MaskFormer [9] and Mask2Former [11] proposes a mask classifier with learnable queries and specialized designs for mask prediction. OneFormer [19] presents a universal image segmentation framework that unifies segmentation with a multi-task train-once design and achieves state-of-the-art performance. Recently, a series of SAM models [20, 40, 22, 21] proposes to build a foundation model for promptable segmentation with strong generalization. Given that the AVS task entails segmentation, these methodologies have significantly contributed to our work.
@@ 2.3 Audio-Visual Segmentation
AVS is an emerging task that aims to locate sounding sources by predicting pixel-wise maps and attracts many researchers [45, 13, 26, 23, 37, 31, 33, 30]. AVSBench [46] first constructs the audio-visual segmentation benchmark and proposes a baseline method that designs a temporal pixel-wise audio-visual interaction module (TPAVI) to inject audio semantics as guidance for the visual segmentation process. AVSegformer [13] proposes a transformer architecture that directly introduces audio features into the transformer decoder, enabling the network to attend to interested visual features selectively. CATR [23] proposes a combinatorial dependence fusion approach that comprehensively accounts for the spatial-temporal dependencies of audio-visual combination. Some methods [15, 30, 31] take advantage of the generative manners with latent diffusion model or variational auto-encoder to address AVS task. In addition, AV-SAM [33], GAVS [37], and BAVS [26] utilize the large foundation model to bootstrap audio-visual segmentation. Different from the above methods [45, 13, 37, 26], our proposed COMBO rethinks AVS from bilateral relations of three entanglements, which enhances the model’s representation ability by exploring the pixel, modality and temporal inherent relationships.
@@ 3.1 Bilateral Visual Features Extraction
Maskige as Prior Knowledge. To solve the aforementioned issues, we believe that a feasible strategy is to incorporate the knowledge of the foundation model into visual features as a pixel entanglement, which is memory-efficient. Specifically, as depicted in the left area of Fig. 2, we introduce a Maskige generator. Given the input image $x\in\mathbb{R}^{3\times H\times W}$ and the frozen foundation model, one can derive the class-agnostic masks $c\in\mathbb{R}^{K\times H\times W}$, where K represents the number of potential targets. It is essential to highlight that K is dynamic and varies with input images. Thus, we first amplify the quantity of class-agnostic masks from $K$ to $N$ with zero masks and obtain a series of binary masks, where N is predetermined. However, given that the output of SAM is a series of binary masks, it is difficult to integrate them into visual features. Consequently, inspired by [6], we introduce Maskige $m\in\mathbb{R}^{3\times H\times W}$, which share the exact dimensions as the input image, to integrate prior knowledge better. To efficiently generate Maskige $m$, we employ a random color encoding function $\mathcal{X}(\cdot):\mathbb{R}^{N}\rightarrow\mathbb{R}^{3}$ that is capable of transforming the binary masks $c\in\mathbb{R}^{N\times H\times W}$ into Maskige $m\in\mathbb{R}^{3\times H\times W}$ without extra training. Specifically, $\mathcal{X}$ is designed to enhance the distinguishability of the Maskige and can be regarded as a linear layer $\mathcal{X}(c)=cA$, where $A\in\mathbb{R}^{N\times 3}$. To facilitate offline inference using the Maskige generator, the value of $A$ is manually set appropriately without additional training. More details are in appendix.
@@
Siam-Encoder Module (SEM). To incorporate the image-like Maskiges as prior knowledge into input frames, we propose intertwining the features of Maskige and visual elements during the feature extraction stage. Accordingly, we design a Siam-Encoder Module, as depicted in Fig. 2. This module encompasses an Image encoder $E_{v}$ and a Maskige encoder $E_{m}$, sharing a common framework. More precisely, for a short video clip with $T$ visual frames $I\in\mathbb{R}^{T\times 3\times H\times W}$, the Maskiges can be generated using the Maskige generator, resulting in $M\in\mathbb{R}^{T\times 3\times H\times W}$. Subsequently, we extract multiple output features from the 2nd-4th stages of both the image encoder and Maskige encoder, respectively. This process can be defined as follows:
@@
in which $F_{v_{i}}$ and $F_{m_{i}}\in\mathbb{R}^{T\times\frac{H}{2^{i+1}}\times\frac{W}{2^{i+1}}\times C_{i}}$ $C_{i}$ represents the dimensions of the $i$th-stage output features. To integrate the Maskige features into COMBO, we introduce channel-weighted blocks that augment the original visual features. The formula can be written as follows:
@@
where GAP stands for global average pooling, and $W\in\mathbb{R}^{C_{i}\times C_{i}}$ represents the linear weight. For simplicity, the bias is omitted in this context. Having obtained the Maskiges as prior information to boost pixel-level entanglement with visual features, the next critical aspect is exploring the modality entanglement between audio and visual signals.
@@ 3.2 Audio-Visual Bilateral Fusion
Audio Feature Extraction. For an audio clip corresponding to the input frames, we adopt VGGish [17] to extract audio features following [45]. Initially, the audio clip is resampled to yield a 16kHz mono output $A_{mono}\in\mathbb{R}^{N_{samples}\times 96\times 64}$, where $N_{samples}$ is related to the duration of the audio. Then, a short-time Fourier transform is performed to yield a mel spectrum, denoted as $A_{mel}\in\mathbb{R}^{T\times 96\times 64}$. Finally, the mel spectrum is subsequently fed into the VGGish model, resulting in the extraction of audio features $F_{a}\in\mathbb{R}^{T\times D}$, where $T$ denotes the number of frames and $D$ represents dimension of the audio.
@@
Bilateral-Fusion Module (BFM). We initially employ the pixel decoder [47] to gradually upsample visual features $F_{v_{i}}$ derived from the SEM to further generate high-resolution per-pixel embeddings $P_{i}\in\mathbb{R}^{T\times\frac{H}{2^{i+1}}\times\frac{W}{2^{i+1}}\times C},\;i\in\{1,2,3,4\}$, where $C$ symbolizes the output channel. Then, we design a Bilateral-Fusion Module (BFM) for constructing a bidirectional audio-visual mapping to assist with segmenting the sounding objects.
As shown in Fig. 3, the BFM utilizes audio features $F_{a}\in\mathbb{R}^{T\times D}$, in conjunction with the largest pixel-level embeddings $P_{1}\in\mathbb{R}^{T\times H_{1}\times W_{1}\times C}$ as inputs which can propagate ample fine-grained semantic information to audio features. Here, $H_{1}=H/4,;W_{1}=W/4$. In order to incorporate both two signals information, bilateral attention is designed within the BFM. Specifically, we initially add fixed sine spatial positional encodings and learnable positional encodings to $P_{1}$ and $F_{a}$, respectively. Next, in order to integrate audio-visual modalities in a memory-efficient way, the SFM comprises four point-wise linear layers that map $P_{1}$ and $F_{a}$ to intermediate representations with dimension $d$. These representations share queries and keys with queries $Q=P_{1}W^{Q}$, keys $K=F_{a}W^{K}$, visual values $V_{v}=P_{1}W_{V}^{v}$, and audio values $V_{a}=P_{1}W_{V}^{a}$. Following the mapping process, the bilateral attention can be computed as follows:
@@
where $d$ is the embedding dim. And $(QK^{T}/\sqrt{d})$ is calculated only once, which is more efficient.
After propagating audio information into fine-grained visual features and injecting visual information into audio features, we proceed by expanding fused audio features $F_{a}$ added with learnable embeddings within the Transformer Decoder as object queries. Additionally, following [10], we generate the output classes $O^{cls}$ by incorporating the per-pixel embeddings $P_{4},P_{3},P_{2}$ into the Transformer Decoder. We acquire the predicted masks $O^{mask}$ by multiplying the output embeddings from the Transformer Decoder with the fused embedding $P_{1}$.

### 07 Apr 2024 v2

@@ Abstract
Recently, an audio-visual segmentation (AVS) task has been introduced, aiming to group pixels with sounding objects within a given video. This task necessitates a first-ever audio-driven pixel-level understanding of the scene, posing significant challenges. In this paper, we propose an innovative audio-visual transformer framework, termed COMBO, an acronym for COoperation of Multi-order Bilateral relatiOns. For the first time, our framework explores three types of bilateral entanglements within AVS: pixel entanglement, modality entanglement, and temporal entanglement. Regarding pixel entanglement, we employ a Siam-Encoder Module (SEM) that leverages prior knowledge to generate more precise visual features from the foundational model. For modality entanglement, we design a Bilateral-Fusion Module (BFM), enabling COMBO to align corresponding visual and auditory signals bi-directionally. As for temporal entanglement, we introduce an innovative adaptive inter-frame consistency loss according to the inherent rules of temporal. Comprehensive experiments and ablation studies on AVSBench-object (84.7 mIoU on S4, 59.2 mIou on MS3) and AVSBench-semantic (42.1 mIoU on AVSS) datasets demonstrate that COMBO surpasses previous state-of-the-art methods. Project page is available at https://yannqi.github.io/AVS-COMBO.
Figure 1: Comparison between the proposed COMBO and existing state-of-the-art methods. Our COMBO is the first work to simultaneously explore multi-order bilateral relations in modality, temporal and pixel levels.
@@ 1 Introduction
Human visual attention is often hear-guided, i.e., we tend to focus on the object with sounds [4]. For example, when we hear a cat meow, we pay more attention to the cat than other objects due to the strong association between the meow and the cat. Inspired by this potential interaction of auditory and visual signals, the cross-modal studies of vision and hearing have attracted numerous researchers, such as the audio-visual correspondence [2, 1], which only aims to match visual images and audio signals to the same scene, and sound source localization [18, 34, 5] which further seeks to locate the vocal visible regions. However, they have only focused on audio-visual tasks at the image or region levels, lacking pixel-level annotations. Recently, AVSBench [45, 46] integrates audio signals into video segmentation, called Audio-Visual Segmentation (AVS), which comprises two benchmarks: 1) AVSBench-object, which includes single source sound segmentation (S4) and multiple sound sources segmentation (MS3); 2) AVSBench-semantic, which further extends audio-visual semantic segmentation (AVSS) based on AVSBench-object.
Given that audio-visual segmentation is a burgeoning field that spans both audio and visual modalities, it presents a non-trivial task. Generally, when performing AVS, a cross-modal segmentation task involving video, there are mainly three challenges: (1) AVS contains audio and visual modalities, thus demanding explicit alignment of sequential audio features to spatial pixel-level activations; (2) AVS involves temporal information where the state of the current frame is dynamically affected by historical frames, therefore, exploring the correlation between adjacent frames is essential; (3) AVS includes an image segmentation task; compared to 1D audio signals, 2D image signals have more redundant information, which is prone to be affected by background noise, thus requiring precise extraction of features from the image. To resolve the issue (1), the prevailing methods [31, 33, 13, 26, 25] employ matrix multiplication and modified cross-attention module to encode pixel-wise audio-visual interaction. Though impressive, these designs ignore the temporal dependence of adjacent frames that have been proven to be important for AVS. To solve this problem, some methods [45, 15, 23] partition temporal relations into consideration to explore both issues (1) and (2) simultaneously since AVS is a cross-modal video task. Nevertheless, their approaches rely too much on implicit inter-frame relations, leading to inaccurate associations. Regarding issue (3), CSMF [3] leverages frozen large foundation models to extract pure visual features for AVS. However, it independently tackles the audio and visual signals by naively combining several existing foundation models, resulting in sub-optimal performance.
@@
In this work, we explore three potential bilateral entanglements: pixel entanglement, modality entanglement, and temporal entanglement. Pixel entanglement refers to the interdependent relationship between an image and its corresponding mask. Since background noise in the image leads to inaccuracies in the image-to-mask prediction process, it is essential to utilize external masks from the foundation models to entangle the input image to assist the model. Therefore, we construct a Siam-Encoder Module (SEM) as a visual feature extractor to facilitate more precise visual features, which can liberate from the constraints of foundation models than other methods [33, 37, 26]. Besides, as for the alignments of the audio and visual signals, we explore the modality entanglement between audio and visual components to amplify the efficiency of cross-modal matching. Contrary to existing single-fusion methods [45, 13], we believe that the cooperation between the two modalities can produce a positive effect. Inspired by [24], we initially propose a potent and memory-efficient bidirectional audio-visual fusion module called Bilateral-Fusion Module (BFM). Our BFM amplifies the spatial awareness of visual features relevant to sounding objects and strengthens the attention of audio signals embodying visual targets. Moreover, the audio-visual tasks contain a solid temporal entanglement. Thus, we design an adaptive inter-frame consistency loss to better harness this inherent characteristic.
@@
  • We show that COMBO significantly outperforms existing state-of-the-art approaches in the challenging AVSBench-object and AVSBench-semantic datasets.
@@ 2 Related Work
2.1 Sound Source Localization
Sound source localization aims to estimate the position of a sound source in a video sequence, which is the most related task to the audio-visual segmentation task. LVS [5] utilizes a hard-mining strategy and a contrastive learning mechanism to discriminate challenging image fragments. DSOL [18] executes class-aware sounding object localization from mixed sound, which initially focuses on learning robust object representations from single-source localization. MSSL [34] localizes multiple sound sources in unconstrained videos without pairwise sound-object annotations. This approach involves the development of a two-stage learning framework, followed by the execution of cross-modal feature alignment. The pioneering methods in these areas have significantly inspired our research on AVS.
@@ 2.2 Semantic Segmentation
Semantic segmentation is a fundamental task that requires pixel-level classification. Early researchers take the Fully Convolutional Networks (FCN) [27] as the dominant approach and focus on aggregating long-range context in the feature map. PSPNet [42] performs spatial pyramid pooling at several grid scales. DeepLab [7, 8] utilizes atrous convolutions with different atrous rates. Furthermore, some methods [38, 35, 43] replace traditional convolutional backbone with transformer-based architectures. MaskFormer [9] and Mask2Former [11] propose a mask classifier with learnable queries and specialized designs for mask prediction. OneFormer [19] presents a universal image segmentation framework that unifies segmentation with a multi-task train-once design. Recently, a series of SAM models [20, 40, 22, 21] propose to build a foundation model for promptable segmentation with strong generalization. Given that the AVS task entails segmentation, these studies have significantly contributed to our work.
@@ 2.3 Audio-Visual Segmentation
AVS is an emerging task that aims to locate sounding sources by predicting pixel-wise maps and attracts many researchers [46, 13, 26, 23, 37, 31, 33, 30]. AVSBench [46] first constructs the audio-visual segmentation benchmark and proposes a temporal pixel-wise audio-visual interaction module (TPAVI) to inject audio semantics as guidance for the visual segmentation process. AVSegformer [13] proposes a transformer architecture that introduces audio features into the transformer decoder, enabling the network to attend to interested visual features selectively. CATR [23] proposes a combinatorial dependence fusion approach that comprehensively accounts for the spatial-temporal dependencies of audio-visual combination. Some methods [15, 30, 31] take advantage of the generative manners with latent diffusion model or variational auto-encoder to address AVS task. In addition, AV-SAM [33], GAVS [37], and BAVS [26] utilize the large foundation model to bootstrap audio-visual segmentation. Different from the above methods [45, 13, 37, 26], our proposed COMBO rethinks AVS from bilateral relations of three entanglements, which enhances the model’s representation ability by exploring the pixel, modality and temporal inherent relationships.
@@ 3.1 Bilateral Visual Features Extraction
Maskige as Prior Knowledge. To solve the aforementioned issues, we believe that a feasible strategy is to incorporate the knowledge of the foundation model into visual features as a pixel entanglement, which is memory-efficient. Specifically, as depicted in the left area of Fig. 2, we introduce a Maskige generator. Given the input image $x\in\mathbb{R}^{3\times H\times W}$ and the frozen foundation model, one can derive the class-agnostic masks $c\in\mathbb{R}^{K\times H\times W}$, where $K$ represents the number of potential targets. It is essential to highlight that $K$ is dynamic and varies with input images. Thus, we first amplify the quantity of class-agnostic masks from $K$ to $N$ with zero masks and obtain a series of binary masks, where $N$ is predetermined. However, given that the output of SAM is a series of binary masks, it is difficult to integrate them into visual features. Consequently, inspired by [6], we introduce Maskige $m\in\mathbb{R}^{3\times H\times W}$, which shares the exact dimensions as the input image, to integrate prior knowledge better. To efficiently generate Maskige $m$, we employ a random color encoding function $\mathcal{X}(\cdot):\mathbb{R}^{N}\rightarrow\mathbb{R}^{3}$ that is capable of transforming the binary masks $c\in\mathbb{R}^{N\times H\times W}$ into Maskige $m\in\mathbb{R}^{3\times H\times W}$ without extra training. Specifically, $\mathcal{X}$ is designed to enhance the distinguishability of the Maskige and can be regarded as a linear layer $\mathcal{X}(c)=cA$, where $A\in\mathbb{R}^{N\times 3}$. To facilitate offline inference using the Maskige generator, the value of $A$ is manually set appropriately without additional training. More details are in Appendix.
@@
Siam-Encoder Module (SEM). To incorporate the image-like Maskiges as prior knowledge into input frames, we propose intertwining the features of Maskige and visual elements during the feature extraction stage. Accordingly, we design a Siam-Encoder Module, as depicted in Fig. 2. This module encompasses an Image encoder $E_{v}$ and a Maskige encoder $E_{m}$, sharing a common framework. More precisely, for a short video clip with $T$ frames $I\in\mathbb{R}^{T\times 3\times H\times W}$, the Maskiges can be generated using the Maskige generator, resulting in $M\in\mathbb{R}^{T\times 3\times H\times W}$. Subsequently, we extract multiple output features from both the image encoder and Maskige encoder, respectively. This process can be defined as follows:
@@
in which $F_{v_{i}}$ and $F_{m_{i}}\in\mathbb{R}^{T\times\frac{H}{2^{i+1}}\times\frac{W}{2^{i+1}}\times C_{i}}$ $C_{i}$ represents the dimension of the $i$-th stage output features. To integrate the Maskige features into COMBO, we introduce channel-weighted blocks that augment the original visual features. The formula can be written as follows:
@@
where GAP stands for global average pooling, and $W\in\mathbb{R}^{C_{i}\times C_{i}}$ represents the linear weight. For simplicity, the bias is omitted in this context. After obtaining the Maskiges as prior information to boost pixel-level entanglement with visual features, the next critical aspect is exploring the modality entanglement between audio and visual signals.
@@ 3.2 Audio-Visual Bilateral Fusion
Audio Feature Extraction. For an audio clip corresponding to the input frames, we adopt VGGish [17] to extract audio features following [45]. Firstly, the audio clip is resampled to yield a 16kHz mono output $A_{mono}\in\mathbb{R}^{N_{samples}\times 96\times 64}$, where $N_{samples}$ is related to the duration of the audio. Then, a short-time Fourier transform is performed to yield a mel spectrum, denoted as $A_{mel}\in\mathbb{R}^{T\times 96\times 64}$. Finally, the mel spectrum is subsequently fed into the VGGish model, resulting in the extraction of audio features $F_{a}\in\mathbb{R}^{T\times D}$, where $T$ denotes the number of frames and $D$ represents dimension of the audio.
@@
Bilateral-Fusion Module (BFM). We initially employ the pixel decoder [47] to gradually upsample visual features $F_{v_{i}}$ derived from the SEM to further generate high-resolution per-pixel embeddings $P_{i}\in\mathbb{R}^{T\times\frac{H}{2^{i+1}}\times\frac{W}{2^{i+1}}\times C},\;i\in\{1,2,3,4\}$, where $C$ denotes the output channel. Then, we design a Bilateral-Fusion Module (BFM) for constructing a bidirectional audio-visual mapping to assist with segmenting the sounding objects.
As shown in Fig. 3, our BFM utilizes audio features $F_{a}\in\mathbb{R}^{T\times D}$, in conjunction with the largest pixel-level embeddings $P_{1}\in\mathbb{R}^{T\times H_{1}\times W_{1}\times C}$ as inputs which can propagate ample fine-grained semantic information to audio features. Here, $H_{1}=H/4,;W_{1}=W/4$. In order to incorporate both two signals, bilateral attention is designed within our BFM. Specifically, we initially add fixed sine spatial positional encodings and learnable positional encodings to $P_{1}$ and $F_{a}$, respectively. Next, in order to integrate audio-visual modalities in a memory-efficient way, our BFM comprises four point-wise linear layers that map $P_{1}$ and $F_{a}$ to intermediate representations with dimension $d$. These representations share queries and keys with queries $Q=P_{1}W_{Q}$, keys $K=F_{a}W_{K}$, visual values $V_{v}=P_{1}W_{V}^{v}$, and audio values $V_{a}=P_{1}W_{V}^{a}$. Following the mapping process, the bilateral attention is as follows:
@@
where $d$ is the embedding dimension. And $(QK^{T}/\sqrt{d})$ is calculated only once, which is more efficient.
After BFM, we proceed by expanding fused audio features $F_{a}$ added with learnable embeddings within the transformer decoder as object queries. Additionally, following [10], we generate the output classes $O^{cls}$ by incorporating the per-pixel embeddings $P_{4},P_{3},P_{2}$ into the transformer decoder. We acquire the predicted masks $O^{mask}$ by multiplying the output embeddings from the transformer decoder with the fused embedding $P_{1}$.